# 1FLAT: a Firmamento-based catalog of AGN in Fermi-LAT high Galactic latitude $\gamma$-ray sources

### v1

@@ Abstract
We present a systematic reassessment of 5,062 high–Galactic latitude gamma-ray sources from the Fermi-LAT 4FGL-DR4 catalog using Firmamento , a web-based platform for multi-frequency source discovery and analysis. Our goal is to provide an independent evaluation of LAT $\gamma$-ray  sources associations through alternative spectral and spatial methods that combine recent and legacy survey data, supplemented by human supervision of spectral energy distributions (SEDs), source morphology, flux variability, and template-based comparisons.
Firmamento confirms the 4FGL-DR4 and 4LAC-DR3 counterparts or unassociated sources in 4,493 cases (88.8%), demonstrating the robustness of both approaches. Beyond this general agreement, we identify 421 new blazar counterparts among previously unassociated sources, thereby reducing the fraction of unidentified extragalactic Fermi-LAT sources from 25% to 17%. In addition, in 64 cases we find alternative blazar associations, while in 49 instances, we do not confirm the 4FGL-DR4 association. For all confirmed blazar counterparts we provide homogeneous estimates of synchrotron peak frequency and peak flux using machine-learning and template-based methods; these agree with 4LAC-DR3 values in most cases, though significant discrepancies appear for a few dozen sources, often due to improved X-ray coverage.
The primary outcome of this work is the First Firmamento  Catalog (1FLAT), made publicly available through the Firmamento platform (https://firmamento.nyuad.nyu.edu), where all related multi-wavelength data and images are available. The project involved extensive manual validation and benefited from the active participation of graduate and undergraduate students, highlighting the platform’s value for both research and education.
@@ 1 Introduction
A small fraction of active galactic nuclei (AGN) launch powerful relativistic jets that emit radiation across the entire electromagnetic spectrum, making them prominent sources in $\gamma$-ray surveys (V. Beckmann & C. Shrader, 2012; C. D. Dermer & G. Menon, 2009). When these jets are aligned with the observer’s line of sight, the resulting relativistic Doppler boosting amplifies the jet’s luminosity, often outshining other emission components of the AGN and leading to their classification as blazars  (P. Padovani et al., 2017). Blazars exhibit a characteristic multi-wavelength emission, with a radio-to-$\gamma$-ray  Spectral Energy Distribution (SED) that typically shows two broad components (A. Abdo et al., 2010). The low-energy component, spanning from radio to optical frequencies and sometimes extending to the X-ray band, is attributed to synchrotron radiation from relativistic electrons spiraling in magnetic fields, while the high-energy emission (X-rays to gamma-rays) arises from inverse Compton scattering of lower-energy photons or other non-thermal processes such as hadronic interactions.
@@
Although these catalogs have significantly advanced our understanding of the $\gamma$-ray sky, a considerable fraction of sources, approximately 25% in the case of 4FGL, remain unidentified due to the lack of clear counterparts at other wavelengths. Among those, S. Abdollahi et al. (2022) report that it is very likely that those at high galactic latitudes are likely unassociated blazars. Identifying the nature of these unidentified sources is crucial for completing the census of $\gamma$-ray emitters and potentially discovering new classes of astrophysical objects.
To address this challenge, we present a systematic reassessment of the AGN counterparts of high-Galactic latitude $\gamma$-ray sources in the 4FGL, utilizing Firmamento (D. Tripathi et al., 2024) (hereafter [Uncaptioned image]), a novel web-based platform specifically developed for the search and identification of multi-frequency counterparts of X-ray and $\gamma$-ray sources. [Uncaptioned image] employs advanced data-handling capabilities, including machine learning models and specialized data science tools (Y. L. Chang et al., 2020; T. Glauch et al., 2022; P. Giommi et al., 2024a). Our search is primarily spectromorphological. We obtain multi-wavelength data through a survey search (see App. A), compare the spatial morphology and the spectral energy distribution relations to match candidates (D. Tripathi et al., 2024).
With respect to LAT, we directly access multi-wavelength survey data, rather than source catalogs, including a number of recent surveys that were not available when 4FGL and 4LAC were prepared. Another important difference is that we include source-by-source human validation on the Spectral Energy Distribution (SED) as well as the multi-wavelength sky maps to supervise the association. Because of this, an accurate false-positive rate cannot be firmly statistically evaluated, although we discuss the robustness of our results throughout the work. Furthermore, in line with M. Ajello et al. (2022) we compute the synchrotron peak and flux at this frequency and compare these values with those of 4LAC. This massive manual work benefited from the active participation of graduate and undergraduate students through [Uncaptioned image]’s user-friendly interface and commitment to educational engagement.
The structure of this paper is as follows. Sec. 2 provides a concise summary of the features of the [Uncaptioned image] platform (presented extensively in (D. Tripathi et al., 2024)) used for this work. Sec. 3 presents the main numerical results of our analysis, discussing both the agreement and discrepancies with previous LAT catalogs and introduce the 1FLAT catalog. In Sec. 4 we study more in depth the properties of our classification, investigating the possible reasons for discrepancies with LAT. In Sec. 5 we discuss the educational engagement utilized in this project. Finally, in Sec. 6 we discuss and summarize our findings.
@@ 2.1 The Firmamento Platform
Further options to investigate source candidates are an Aladin-based multi-wavelength skymap displaying the retrieved catalogs data, benchmark SEDs (LSP, ISP, HBL) and host galaxy and blue bump templates to superimpose to data for source validation.
@@ 2.2 Workflow for 4FGL Sources
We conducted an independent search for blazars among the high-Galactic latitude ($|b|>10^{\circ}$) $\gamma$-ray sources of the LAT 4FGL catalog (S. Abdollahi et al., 2022), with a starting dataset of 5,062 sources. We prepared an ASCII file with 4FGL Source name Source_Name, Right Ascension, Declination RAJ2000, DEJ2000, Long and Short radii of error ellipse at 95% confidence Conf_95_SemiMajor, Conf_95_SemiMinor, Position angle (eastward) of the long axis from celestial North Conf_95_PosAng and the Name of identified or likely associated source name and class ASSOC1, CLASS along with its position RA_Counterpart, DEC_Counterpart. From the 4LAC we also took the ASSOC1, CLASS1 fields. We also annonated the 4FGL association probabilities ASSOC_PROB_BAY, ASSOC_PROB_LR. We remark that we did not consider the 4FGL secondary association ASSOC2, CLASS2 because no position is given. However, we took note of those secondary association for discussion of some specific cases. This type of information was input in [Uncaptioned image] which features a special mode for user-input table-data. We run the ERCI tool on all sources one by one. ERCI can provide none, single or multiple candidate counterparts. Whenever ERCI identified one or more candidates, we verified the validity of each potential counterpart based on its multi-frequency morphological properties (via the Aladin skymaps) and its SED. Particular attention was given to the SED to ensure the synchrotron peak was consistent with the $\gamma$-ray intensity and spectral slope reported in 4FGL, thus confirming or rejecting its association as a blazar. Cases with multiple plausible candidates were resolved by selecting the counterpart with the most compelling SED. While specific quantitative thresholds for all parameters are complex and depend on the multi-dimensional parameter space explored by ERCI, the final selection of counterparts involved a careful visual inspection of the SEDs by experienced users and, in many cases, by undergraduate and graduate students under expert supervision (see Sec. 5). The synchrotron peak position estimated by BLAST and W-Peak (T. Glauch et al., 2022; P. Giommi et al., 2024a) provided additional validation of the blazar nature of the counterparts.
The coincidence with the proposed 4FGL or [^1] There were cases in which LAT sources were associated to blazar in both 4FGL and 4LAC but some only in 4FGL and not in 4LAC and vice-versa. We required a blazar classification in either of the two catalogs. 4LAC candidates was also finally scrutinized based on the position (and naming) of the LAT proposed association.
The step-by-step procedure is reported in App. B.
@@ 3.1 Numerical Overview
The results of the search for counterparts to the 5,062 4FGL $\gamma$-ray sources above the Galactic plane ($|b|>10^{\circ}$) with [Uncaptioned image]  are discussed in Sec. 2 and are summarized in Tab. 1.
@@
$\triangleright$ [Uncaptioned image] does not confirm LAT association 16 0.3
@@
Table 1: Summary of the classification of 4FGL/4LAC sources made with [Uncaptioned image]. Percentages are calculated with respect to the total number of 4FGL sources, excluding those associated with Galactic sources in the catalog. The upper block lists the cases where there is agreement with LAT, while the lower block gives the cases where there are disagreements.
@@
[Uncaptioned image]  is in disagreement with 4FGL or 4LAC for 569 (11.2%) $\gamma$-ray  sources. Among these, the largest population is that of new blazar associations discovered with [Uncaptioned image] , which amounts to 421 (8.3%) objects. In 64 cases (1.4%) we instead find a different, and more plausible, blazar counterpart. In 49 cases (1.0%) we cannot confirm a candidate that instead is claimed by 4FGL. In 16 cases (0.3%), although we observe the presence of a possible counterpart, its nature cannot be confirmed with high confidence. Finally, [Uncaptioned image] finds different counterparts for 19 (0.4%) sources, including 18 galaxies and 1 is a Galactic object. In the following sections, we discuss individual cases, focusing on blazars, the class of sources most relevant to this work.
@@ 3.2 Agreement between Firmamento and 4FGL
We focus here on the 4,493 (88.8%) sources in which there is an agreement between [Uncaptioned image] and 4FGL.
@@ 3.2.1 Confirmed blazars
Figure 1: The most common situation exemplified by the case of 4FGL J2221.8+3358 where Refer to caption  and 4FGL/4LAC identify the same $\gamma$-ray  blazar. Left: Refer to caption proposes a single counterpart coincident with the LAT 4LAC counterpart (yellow circle). The purple ellipse in the figure represents the LAT error region. Right: The SED of CRATES J222152+3335844, the counterpart of 4FGL J2221.8+3358.
In Tab. 2 we report the blazar classification obtained with [Uncaptioned image] for three CLASS values: bll,fsrq,bcu. One can see that more than half of the 4FGL bll are HSP, basically all the [Uncaptioned image] LSP are FSRQs while we find that half of the 4FGL bcus are LSP. We did not report the statistics for the 4FGL blazars with different classification as their contribution is minor.
@@
The estimation of the synchrotron peak, not present in 4FGL, was instead provided in 4LAC. We report the comparison of our classification with that of 4LAC in Sec. 4.2 over our entire blazar sample.
@@ 3.2.2 Confirmed missing counterparts
For 854 4FGL sources (16.9%), we concur with LAT on the absence of a plausible counterpart. This is likely due to the lack of availability of multi-wavelength data at least in one of the bands that characterize a blazar (e.g. X-ray or optical/infrared). In such cases, the [Uncaptioned image] algorithm does not reach the threshold for plausible counterpart proposal and return a null association. We remark that if future data will be added, the same algorithm could be run again to re-evaluate the association.
We remark that among these entries, 14 sources have a non-empty 4FGL_CLASS2, with the following classifications: agn: 8, unk: 5, nlsy1: 1.
3.2.3 Confirmed non-AGN counterparts
@@ 3.6 Non-Confirmation of 4FGL/4LAC Associations
In 49 other cases (1.0%), we consider the 4FGL or 4LAC counterparts as unreliable and no alternative counterparts were found by ERCI within the error region. These source were mostly bcus in 4FGL_CLASS1.
@@ 3.8 The 1FLAT  catalog
Our findings are compiled in a catalog called 1FLAT  (first Firmamento  LAT AGN Table). The catalog is primarily focused on blazars, although we report some additional results to enable interested scientist to further explore our work. We instead do not list Galactic objects or radio galaxies. In summary the 1FLAT  includes:
  • Blazars: 3,456 blazars found by [Uncaptioned image] in agreement with LAT  421 new [Uncaptioned image] blazar associations; 64 blazars for which [Uncaptioned image] finds a different blazar candidate than LAT
@@
  • Unassociated: 854 $\gamma$-ray  sources for which both [Uncaptioned image] and LAT do not find an association and 49 sources for which [Uncaptioned image] does not find a valid association whereas LAT does
  • Galaxies: 20 galaxies found by [Uncaptioned image] in agreement with 4FGL and 18 galaxies for which [Uncaptioned image] finds a different blazar candidate than LAT
@@ 3.9 Comparison with Other Association Methods
The identification of 421 new blazar associations among previously unassociated $\gamma$-ray  sources, along with the association of 64 sources with counterparts that are different from those reported in 4FGL, represent two major outcomes of this work and highlight important differences with respect to the 4FGL catalog results. In the following, we briefly compare our [Uncaptioned image] -based technique with the Bayesian and Likelihood Ratio methods adopted by the LAT collaboration.
The 4LAC and [Uncaptioned image] methods are fundamentally different and depend on markedly different amounts of information. As detailed in Sec. 2.1, the algorithm implemented in [Uncaptioned image] combines very large amount of multi-frequency data with information on source variability and spatial extension. [Uncaptioned image] determines whether a given error region includes one or more sources that are likely to be blazars or other types of multi-frequency emitters through a two-step process: first it analyzes the shape of the broadband SED of all radio and X-ray sources in the requested area, constructed from approximately 50 catalogs and survey data; then, it builds a more detailed SED from 90 catalogs and spectral databases (see App. A) , which is visually inspected by our team to confirm or reject the candidate(s). This final human intervention will eventually be replaced by a machine learning tool trained on the results of this and similar works.
@@
Since the publication of the first Fermi-LAT catalogs, a number of independent teams have also devised alternative methods to aid in the identification of unassociated $\gamma$-ray  sources. Some of these approaches use machine learning techniques that rely solely on $\gamma$-ray  data (D. Salvetti et al., 2017), while others combine $\gamma$-ray  and X-ray data (A. Kaur et al., 2019).
To support the identification process, R. D’Abrusco et al. (2019) selected a large sample of blazar candidates based on radio data and infrared colors, while Y.-L. Chang et al. (2019) compiled a sample of high-energy-peaked blazars, many of which were expected to be detected by Fermi-LAT and are indeed listed among the confirmed and newly identified blazars in the 1FLAT catalog. More recent works have focused on the selection of samples of blazars that are detected in the very-high-energy (VHE) band (e.g. B. Arsioli et al., 2025; A. Neronov & D. Semikoz, 2025). A detailed comparison between our work and these efforts — some of which are considered in the 4FGL-DR4 and 4LAC-DR3 papers — is beyond the scope of this study.

### v2

@@ Abstract
We present a systematic reassessment of 5,062 high–Galactic latitude gamma-ray sources from the Fermi-LAT 4FGL-DR4 catalog using Firmamento , a web-based platform for multi-frequency source discovery and analysis. Our goal is to provide an independent evaluation of LAT $\gamma$-ray  source associations using alternative spectral and spatial methods that integrate both recent and legacy survey data. The evaluation is further refined through human supervision of SEDs, source morphology, flux variability, and template-based comparisons. Firmamento confirms the 4FGL-DR4 and 4LAC-DR3 counterparts or unassociated sources in 4,493 cases (88.8%), demonstrating the robustness of both approaches. Beyond this general agreement, we identify 421 new blazar counterparts among previously unassociated sources, thereby reducing the fraction of unidentified extragalactic Fermi-LAT sources from 25% to 17%. In addition, in 64 cases we find alternative blazar associations, while in 49 instances, we do not confirm the 4FGL-DR4 association. For all confirmed blazar counterparts we provide homogeneous estimates of synchrotron peak frequency and peak flux using machine-learning and template-based methods. The results agree with 4LAC-DR3 values in most cases, though significant discrepancies appear for a few dozen sources, often due to improved X-ray coverage. The primary outcome of this work is the 1st Firmamento LAT AGN Table (1FLAT), made publicly available through the Firmamento platform (https://firmamento.nyuad.nyu.edu), where all related multi-wavelength data and images are available. The project involved extensive manual validation and benefited from the active participation of graduate and undergraduate students, highlighting the platform’s value for both research and education.
@@ 1 Introduction
A small fraction of active galactic nuclei (AGN) launch powerful relativistic jets that emit radiation across the entire electromagnetic spectrum, making them prominent sources in $\gamma$-ray surveys (V. Beckmann & C. Shrader, 2012; C. D. Dermer & G. Menon, 2009). When these jets are aligned with the observer’s line of sight, the resulting relativistic Doppler boosting amplifies the jet’s luminosity, often outshining other emission components of the AGN and leading to their classification as blazars  (P. Padovani et al., 2017). Blazars exhibit a characteristic multi-wavelength emission, with a radio-to-$\gamma$-ray  Spectral Energy Distribution (SED) that typically shows two broad components (A. Abdo et al., 2010). The low-energy component, spanning from radio to optical frequencies, and sometimes extending to the X-ray band, is attributed to synchrotron radiation from relativistic electrons spiraling in magnetic fields. The high-energy emission (X-rays to gamma-rays) arises from inverse Compton scattering of lower-energy photons or from other non-thermal processes such as hadronic interactions.
@@
Although these catalogs have significantly advanced our understanding of the $\gamma$-ray sky, a considerable fraction of sources, approximately 25% in the case of 4FGL, remain unidentified due to the lack of clear counterparts at other wavelengths. Among those, S. Abdollahi et al. (2022) report that it is very likely that those at high Galactic latitudes are likely unassociated blazars. Identifying the nature of these sources is crucial for completing the census of $\gamma$-ray emitters and potentially discovering new classes of astrophysical objects.
To address this challenge, we present a systematic reassessment of the AGN counterparts of high-Galactic latitude $\gamma$-ray sources in the 4FGL, utilizing Firmamento (D. Tripathi et al., 2024) (hereafter [Uncaptioned image]), a novel web-based platform specifically developed for the search and identification of multi-frequency counterparts of X-ray and $\gamma$-ray sources. [Uncaptioned image] employs advanced data-handling capabilities, including machine learning models and specialized data science tools (Y. L. Chang et al., 2020; T. Glauch et al., 2022; P. Giommi et al., 2024a). Our search is primarily spectromorphological. We obtain multi-wavelength data through a survey search (see App. A), compare the spatial morphology and the SED relations to match candidates (D. Tripathi et al., 2024).
With respect to 4FGL, we directly access multi-wavelength survey data, rather than source catalogs, including a number of recent surveys that were not available when 4FGL and 4LAC were prepared. Another important difference is that we include source-by-source human validation on the SED as well as the multi-wavelength sky maps to supervise the association. Because of this, an accurate false-positive rate cannot be firmly statistically evaluated, although we discuss the robustness of our results throughout the work. Furthermore, in line with M. Ajello et al. (2022) we compute the synchrotron peak and flux at this frequency and compare these values with those of 4LAC. This massive manual work benefited from the active participation of graduate and undergraduate students through [Uncaptioned image]’s user-friendly interface and commitment to educational engagement.
The structure of this paper is as follows. Sec. 2 provides a concise summary of the features of the [Uncaptioned image] platform (presented extensively in (D. Tripathi et al., 2024)) used for this work. Sec. 3 presents the main numerical results of our analysis, discussing both the agreement and discrepancies with previous LAT catalogs and introduce the 1FLAT  catalog. In Sec. 4 we study more in depth the properties of our classification, investigating the possible reasons for discrepancies with 4FGL. In Sec. 5 we discuss the educational engagement utilized in this project. Finally, in Sec. 6 we discuss and summarize our findings.
@@ 2.1 The Firmamento Platform
Additional tools for investigating source candidates include an Aladin-based multi-wavelength sky map displaying the retrieved catalog data, as well as benchmark SEDs (LSP, ISP, HSP) and host-galaxy or blue-bump templates that can be superimposed on the data for source validation.
@@ 2.2 Workflow for 4FGL Sources
We conducted an independent search for blazars among the high-Galactic-latitude ($|b|>10^{\circ}$) $\gamma$-ray sources of the LAT 4FGL catalog (S. Abdollahi et al., 2022), with a starting dataset of 5,062 sources. We prepared an ASCII file with 4FGL Source name Source_Name, Right Ascension, Declination RAJ2000, DEJ2000, Long and Short radii of error ellipse at 95% confidence Conf_95_SemiMajor, Conf_95_SemiMinor, Position angle (eastward) of the long axis from celestial North Conf_95_PosAng and the Name of identified or likely associated source name and class ASSOC1, CLASS along with its position RA_Counterpart, DEC_Counterpart. From the 4LAC catalog we took the ASSOC1, CLASS1 fields. We also annonated the 4FGL association probabilities ASSOC_PROB_BAY, ASSOC_PROB_LR. We remark that we did not consider the 4FGL secondary association ASSOC2, CLASS2 because no position is given. However, we took note of those secondary association for discussion of some specific cases. This type of information was entered into [Uncaptioned image] , which features a special mode for user-input table-data. We run the ERCI tool on all sources one by one. ERCI can provide none, single or multiple candidate counterparts. Whenever ERCI identified one or more candidates, we verified the validity of each potential counterpart based on its multi-frequency morphological properties (via the Aladin skymaps) and its SED. Particular attention was given to the SED to ensure the synchrotron peak was consistent with the $\gamma$-ray intensity and spectral slope reported in 4FGL. Cases with multiple plausible candidates were resolved by selecting the counterpart with the most compelling SED. While specific quantitative thresholds for all parameters are complex and depend on the multi-dimensional parameter space explored by ERCI, the final selection of counterparts involved a careful visual inspection of the SEDs by experienced users and, in many cases, by undergraduate and graduate students under expert supervision (see Sec. 5). The synchrotron peak position estimated by BLAST and W-Peak (T. Glauch et al., 2022; P. Giommi et al., 2024a) provided additional validation of the blazar nature of the counterparts.
The coincidence with the proposed 4FGL or [^1] There were cases in which LAT sources were associated to blazars in both 4FGL and 4LAC but some only in 4FGL and not in 4LAC and vice-versa. We required a blazar classification in either of the two catalogs. 4LAC candidates were also finally scrutinized based on the position (and naming) of the proposed association. The step-by-step procedure is reported in App. B.
@@ 3.1 Numerical Overview
The results of the search for counterparts to the 5,062 4FGL $\gamma$-ray sources above the Galactic plane ($|b|>10^{\circ}$) with [Uncaptioned image] , are discussed below and are summarized in Tab. 1.
@@
$\triangleright$ [Uncaptioned image] does not confirm the association 16 0.3
@@
Table 1: Summary of the classification of 4FGL/4LAC sources made with [Uncaptioned image]. Percentages are calculated with respect to the total number of 4FGL sources, excluding those associated with Galactic sources in the catalog. The upper block lists the cases where there is agreement with 4FGL, while the lower block gives the cases where there are disagreements.
@@
[Uncaptioned image]  is in disagreement with 4FGL or 4LAC for 569 (11.2%) $\gamma$-ray  sources. Among these, the largest population is that of new blazar associations discovered with [Uncaptioned image] , which amounts to 421 (8.3%) objects. In 64 cases (1.4%) we instead find a different, and more plausible, blazar counterpart. In 49 cases (1.0%) we cannot confirm a candidate that instead is claimed by 4FGL. In 16 cases (0.3%), although we observe the presence of a possible counterpart, its nature cannot be confirmed with high confidence. Finally, [Uncaptioned image] finds different counterparts for 19 (0.4%) sources, including 18 galaxies and 1 Galactic object. In the following sections, we discuss individual cases, focusing on blazars, the class of sources most relevant to this work.
@@ 3.2 Agreement between Firmamento and 4FGL
We focus here on the 4,493 (88.8%) sources for which there is an agreement between [Uncaptioned image]  and 4FGL.
@@ 3.2.1 Confirmed blazars
Figure 1: The most common situation exemplified by the case of 4FGL J2221.8+3358 where Refer to caption  and 4FGL/4LAC identify the same $\gamma$-ray  blazar. Left: Refer to caption proposes a single counterpart coincident with the 4LAC counterpart (yellow circle). The purple ellipse in the figure represents the 4FGL error region. Right: The SED of CRATES J222152+3335844, the counterpart of 4FGL J2221.8+3358.
In Tab. 2 we report the blazar classification obtained with [Uncaptioned image] for three CLASS values: bll,fsrq,bcu. One can see that more than half of the 4FGL bll are HSPs, almost all the FSRQs are LSPs while we find that half of the 4FGL bcus are LSPs. We did not report the statistics for the 4FGL blazars with different classification as their contribution is minor.
@@
The estimation of the synchrotron peak, not present in 4FGL, is instead provided in 4LAC. We report the comparison of our classification with that of 4LAC in Sec. 4.2 over our entire blazar sample.
@@ 3.2.2 Confirmed missing counterparts
For 854 4FGL sources (16.9%), we concur with 4FGL on the absence of a plausible AGN counterpart. This is partly due to the lack of availability of multi-wavelength data at least in one of the bands that characterize a blazar (e.g. X-ray or optical/infrared). In such cases, the [Uncaptioned image] algorithm does not reach the threshold for plausible counterpart proposal and return a null association. We remark that if future data will be added, the same algorithm could be run again to re-evaluate the association. We remark that among these entries, 14 sources have a non-empty 4FGL_CLASS2, with the following classifications: agn: 8, unk: 5, nlsy1: 1.
3.2.3 Confirmed non-blazar counterparts
@@ 3.6 Non-Confirmation of 4FGL/4LAC Associations
In 49 other cases (1.0%), the 4FGL or 4LAC counterparts are considered unreliable, and ERCI did not find alternative counterparts within the error region. These source were mostly bcus in 4FGL_CLASS1.
@@ 3.8 The 1FLAT  catalog
Our findings are compiled in a catalog called 1FLAT  (first Firmamento  LAT AGN Table). The catalog is primarily focused on blazars, although we report some additional results to enable interested scientist to further explore our work. We instead do not list Galactic objects or non-active galaxies. In summary the 1FLAT  includes:
  • Blazars: 3,456 blazars found by [Uncaptioned image] in agreement with 4FGL; 421 new [Uncaptioned image] blazar associations; 64 blazars for which [Uncaptioned image] finds a different blazar candidate than 4FGL
@@
  • Unassociated: 854 $\gamma$-ray  sources for which both [Uncaptioned image] and 4FGL do not find an association and 49 sources for which [Uncaptioned image] does not find a valid association whereas 4FGL does
  • Radio galaxies: 20 galaxies found by [Uncaptioned image] in agreement with 4FGL and 18 galaxies for which [Uncaptioned image] finds a different blazar candidate than 4FGL
@@ 3.9 Comparison with Other Association Methods
The identification of 421 new blazar associations among previously unassociated $\gamma$-ray  sources represents one of the major outcomes of this work. In addition, we associate 64 sources with counterparts that differ from those reported in 4FGL, highlighting important differences with respect to the 4FGL catalog results. In the following, we briefly compare our [Uncaptioned image] -based technique with the Bayesian and Likelihood Ratio methods adopted by the LAT collaboration.
The 4LAC and [Uncaptioned image] methods are fundamentally different and depend on markedly different amounts of information. As detailed in Sec. 2.1, the algorithm implemented in [Uncaptioned image] combines very large amount of multi-frequency data with information on source variability and spatial extension. [Uncaptioned image] determines whether a given error region includes one or more sources that are likely to be blazars or other types of multi-frequency emitters through a two-step process: first it analyzes the shape of the broadband SED of all radio and X-ray sources in the requested area, constructed from approximately 50 catalogs and survey data; then, it builds a more detailed SED from approximately 90 catalogs and spectral databases (see App. A), which is visually inspected by our team to confirm or reject the candidate(s). This final human intervention will eventually be replaced by a machine learning tool trained on the results of this and similar works.
@@
Since the publication of the first LATcatalogs, a number of independent teams have also devised alternative methods to aid in the identification of unassociated $\gamma$-ray  sources. Some of these approaches use machine learning techniques that rely solely on $\gamma$-ray  data (D. Salvetti et al., 2017), while others combine $\gamma$-ray  and X-ray data (A. Kaur et al., 2019).
To support the identification process, R. D’Abrusco et al. (2019) selected a large sample of blazar candidates based on radio data and infrared colors, while Y.-L. Chang et al. (2019) compiled a sample of high-energy-peaked blazars, many of which were expected to be detected by LATand are indeed listed among the confirmed and newly identified blazars in the 1FLAT  catalog. More recent works have focused on the selection of samples of blazars that are detected in the very-high-energy (VHE) band (e.g. B. Arsioli et al., 2025; A. Neronov & D. Semikoz, 2025). A detailed comparison between our work and these efforts — some of which are considered in the 4FGL-DR4 and 4LAC-DR3 papers — is beyond the scope of this study.